OpenStudy (anonymous):

What value of x makes this proportion true? x/6=36/24

OpenStudy (texaschic101):

x/6 = 36/24.....cross multiply (24)(x) = (36)(6) 24x = 216 -- now divide both sides by 24 x = 216/24 x = 9 <=== your answer just to let you know, proportions are nothing but equivalent fractions. 9/6 = 3/2 36/24 = 3/2

OpenStudy (anonymous):

9. because 36/24 reduces down to 3/2. x/6=3/2 is 9. 9/6. boom